# 3DGraphLLM: Combining Semantic Graphs and Large Language Models for 3D Scene Understanding

### v2

@@ Abstract
A 3D scene graph represents a compact scene model, storing information about the objects and the semantic relationships between them, making its use promising for robotic tasks. When interacting with a user, an embodied intelligent agent should be capable of responding to various queries about the scene formulated in natural language. Large Language Models (LLMs) are beneficial solutions for user-robot interaction due to their natural language understanding and reasoning abilities. Recent methods for creating learnable representations of 3D scenes have demonstrated the potential to improve the quality of LLMs responses by adapting to the 3D world. However, the existing methods do not explicitly utilize information about the semantic relationships between objects, limiting themselves to information about their coordinates. In this work, we propose a method 3DGraphLLM for constructing a learnable representation of a 3D scene graph. The learnable representation is used as input for LLMs to perform 3D vision-language tasks. In our experiments on popular ScanRefer, RIORefer, Multi3DRefer, ScanQA, Sqa3D, and Scan2cap datasets, we demonstrate the advantage of this approach over baseline methods that do not use information about the semantic relationships between objects. The code is publicly available at https://github.com/CognitiveAISystems/3DGraphLLM.
@@ 1 Introduction
In this paper, we consider scene understanding in the context of solving 3D vision-language problems: 3D referred object grounding task, 3D dense scene captioning and 3D visual question answering. The 3D referred object grounding task involves identifying a region in a 3D scene that corresponds to a complex natural language query. This query may describe object properties (e.g., color, size) and spatial relationships (e.g., a mug on a table). A common approach to solving this problem is to assume that one is given a 3D reconstruction of the scene (e.g., a point cloud, mesh, or NeRF). The goal is to predict the bounding boxes of the region that matches the query. The goal of the dense scene captioning task is to describe the selected object. Finally, the goal of the 3D visual question answering task is to generate text answer to various questions about the properties of the scene. It seems promising to explicitly use a three-dimensional scene graph to solve these tasks.
The 3D scene graph not only allows storing multimodal information about individual objects within a scene but also captures the semantic relationships (Wang et al., 2023a; Koch et al., 2024) and hierarchical organization between them (Werby et al., 2024; Honerkamp et al., 2024). Additionally, the graph scene representation enables real-time updates for dynamic environments (Rosinol et al., 2021; Özsoy et al., 2023), and supports the application of graph algorithms for tasks such as navigation (Zhou et al., 2023a; He and Zhou, 2024; Honerkamp et al., 2024) or object search based on textual queries (Feng et al., 2021; Chang et al., 2023; Werby et al., 2024; Gu et al., 2024).
The solving of 3D vision-language tasks (Chen et al., 2020, 2021; Azuma et al., 2022) is crucial for embodied intelligent agents. To interact with the user, an intelligent agent must be able to describe the environment and answer questions about its properties using natural language. Large language models (LLMs) are particularly well-suited for this task, as their advanced capabilities in natural language understanding and common-sense reasoning make them highly effective in interpreting and matching user queries to objects in a scene (Hong et al., 2023a; Wang et al., 2024; Gu et al., 2024). Using LLMs makes it easier to adapt the method to new categories of objects and relationships found in referring expressions. LLMs can also handle complex queries that don’t explicitly mention the class name, but instead describe its function (e.g. "somewhere to sit").
@@
Figure 1: Proposed 3DGraphLLM approach leverages 3D semantic scene graph learnable representation supplied as input to an LLM to perform various 3D vision-language tasks.
The 3D scene description input for LLMs can be represented either as text (Gu et al., 2024; Linok et al., 2024; Werby et al., 2024; Honerkamp et al., 2024; Yang et al., 2024; Yuan et al., 2024), or through learnable representations (Hong et al., 2023a; Chen et al., 2023; Huang et al., 2023; Chen et al., 2024; Cheng et al., 2024), which encode objects and their relationships using significantly fewer tokens and their corresponding embeddings than a textual description of the scene. These learnable representations enhance the performance of the LLM in generating responses to user queries, while also improving response accuracy through adaptation to 3D scenes. However, current methods (Hong et al., 2023a; Chen et al., 2023; Huang et al., 2023; Chen et al., 2024) for 3D vision-language tasks using LLM and learnable 3D scene representations fail to leverage semantic relationships between objects, relying solely on their spatial coordinates.
@@
In this paper, we propose a novel learnable representation of a 3D scene graph called 3DGraphLLM, designed for use as input to a LLM (see Figure 1). This representation consists of a list of learnable embeddings for objects within the scene, where each object is represented by a subgraph containing the object itself along with several of its nearest neighbors. These object subgraphs are provided to the LLM as a sequence of triplets (object1, relation, object2). Semantic relations between objects are embedded using features derived from the semantic edges of the graph, which is generated using state-of-the-art methods for 3D semantic graph generation such as VL-SAT (Wang et al., 2023a). Our experiments demonstrate that incorporating semantic relationships between objects significantly improves the accuracy of LLM responses for 3D vision-language tasks, outperforming baseline approaches for creating learnable scene representations.
@@
  • We introduce 3DGraphLLM, the first method to create a learnable 3D scene graph representation for LLMs, enabling the mapping of semantic relationships between objects in the scene to LLM’s token embedding space.
  • We propose an algorithm that produces a flat sequence of graph embedding tokens using k-nearest neighbor selection with a minimum distance filter between objects, optimizing inference speed by reducing the number of tokens required to describe the scene.
  • 3DGraphLLM shows state-of-the-art results for the 3D referred object grounding task on the Multi3DRefer (Zhang et al., 2023) (+5.8% F1@0.5) and ScanRefer (Chen et al., 2020) (+4.4% Acc@0.5) benchmarks and also for the 3D scene captioning on the Scan2Cap dataset Chen et al. (2021) (CIDEr@0.5 +5.8%).
@@ 2 Related works
Scene Graphs. The concept of a scene graph was initially developed for 2D images, providing a structured representation of a scene’s semantics by incorporating relationships between the semantic elements (Johnson et al., 2015). In the context of images, scene graphs have proven effective for tasks such as content-based image retrieval (Johnson et al., 2015; Pei et al., 2023), 2D referring expression comprehension (Yang et al., 2019a; Shi et al., 2023; Han et al., 2024), image caption (Yang et al., 2019b; Phueaksri et al., 2023), image generation (Johnson et al., 2018; Farshad et al., 2023).
In 3D scenes, a scene graph is commonly used to address robotics challenges such as planning (Werby et al., 2024; Honerkamp et al., 2024), object grounding for navigation (Werby et al., 2024; Gu et al., 2024; Linok et al., 2024; Honerkamp et al., 2024) and manipulation (Honerkamp et al., 2024), as well as scene generation (Zhai et al., 2024; Gao et al., 2024). Our approach is part of a class of methods that utilize an implicit representation of the scene graph, such as OVSG (Chang et al., 2023), which frames the problem of 3D object grounding as subgraph retrieval. 3DGraphQA (Wu et al., 2024) proposes to use the bilinear graph neural network for feature fusion between scene and question graphs for question answering task.  Feng et al. (2021) build a graph based on a text query, which is used to refine the visual graph in order to select from its vertices the one that best fits the description. However, the application scope of this method is limited to specific tasks as 3D referred object grounding with one referred object or question answering. In contrast, we propose a more versatile method capable of solving various 3D vision-language tasks.
@@
3D Language Scene Understanding. 3D scene understanding is a complex computer vision task that involves identifying the semantic, physical, and functional properties of objects, as well as their mutual relations. One of the goals of 3D scene understanding is to develop methods capable of responding to natural language queries about the scene. The queries may correspond to different visual-language tasks such as 3D referred object grounding (Chen et al., 2020; Zhang et al., 2023; Miyanishi et al., 2024), question answering (Azuma et al., 2022), and dense scene captioning (Chen et al., 2021). Recent approaches address these queries by reconstructing the scene as a 3D mesh (Peng et al., 2023) or point cloud (Zhao et al., 2021; Chen et al., 2022; Zhu et al., 2023), often enhanced with instance segmentation (Zhu et al., 2023).
@@
The emergence of transformer models (Vaswani, 2017) has enabled the development of neural network models that create a learnable representation of a scene for answering various language queries. MultiCLIP (Delitzas et al., 2023) proposes to align 3D scene representation with text queries and multi-view 2D CLIP (Radford et al., 2021) embeddings to improve the quality of question answering. 3DVG-Transformer (Zhao et al., 2021) and Vil3DRef (Chen et al., 2022) methods introduce modules for modeling spatial relationships between objects to improve the quality of object grounding. 3D-VisTA (Zhu et al., 2023) presents a transformer model for aligning 3D object and text representations, coupled with an unsupervised pre-training scheme to solve various 3D vision-text problems using specialized task-specific heads. However, these fully supervised approaches face challenges in generalizing to new tasks and domains. In contrast, leveraging large language models (LLMs) for scene understanding enhances generalization capabilities and taps into the extensive knowledge LLMs contain about the physical world (Hong et al., 2023a).
@@
Large Language Models for Scene Understanding. Large language models (LLMs) offer several advantages for scene understanding, notably enhancing the ability to address complex queries that require common knowledge. LLMs can serve as agents that decompose user queries into elementary tasks, which can then be addressed by other methods (Yang et al., 2024; Yuan et al., 2024). Additionally, LLMs can act as an interface for reasoning by processing textual descriptions of the scene as input (Linok et al., 2024; Gu et al., 2024). BBQ (Linok et al., 2024) and ConceptGraphs (Gu et al., 2024) demonstrate that using a text-based graph representation with an LLM interface significantly improves the quality of object retrieval compared to using CLIP features of objects. HOV-SG (Werby et al., 2024) construct a hierarchical graph consisting of objects, rooms, and floors, and demonstrate the effectiveness of such a representation for the task of object grounding given a query containing object location hints. The authors of the MOMA (Honerkamp et al., 2024) method propose using a hierarchical scene graph together with a navigational Voronoi graph as input to LLM to predict a high-level policy for object search for navigation and manipulation. However, using text to describe an object in a scene graph inevitably leads to the loss of some of the information contained in its RGB point cloud. Additionaly, in the case of using a text graph, several hundred tokens may be required to describe one object (its semantic class, pose), which will significantly slow down LLM inference in the case of a large number of objects in the scene.
@@
Recent advancements have successfully integrated point cloud data into LLMs by employing pre-trained point cloud encoders and training adapters to align the resulting representations with the LLM embedding space. 3D-LLM (Hong et al., 2023b) aggregates 3D point cloud features from a sequence of 2D images and then solves the grounding problem as a prediction of a sequence of location tokens added to the LLM dictionary. Chat3D-v2 (Huang et al., 2023) generates a 3D feature for each object in the scene and then treats the grounding problem as an object selection problem. LLA3D (Chen et al., 2023) proposes to use a set of trainable fixed-length query tokens obtained by interacting potential visual cues, text cues, and object point cloud features in a transformer model. Grounded 3D-LLM (Chen et al., 2024) uses referent tokens to decode object masks in point clouds. Additionally, research has demonstrated that incorporating spatial information, such as object coordinates (Huang et al., 2023) or depth maps (Cheng et al., 2024), enhances the accuracy of responses to user queries.
@@ 3 Method
Our approach uses a set of point clouds of scene objects as input. The objects’ point clouds can be obtained either from ground-truth annotations or through state-of-the-art point cloud instance segmentation methods. These point clouds are used to extract scene graph features (see Section 3.1). A scene graph consists of nodes representing the objects and edges corresponding to semantic relationships between them. To convert the scene graph into a token sequence, we represent each object by an identifier, followed by a subgraph comprising the object’s k nearest neighbors. The relationships between an object and its neighbors are encoded as triplets $(object_{i},relation_{ij},object_{j})$. The scheme of the 3DGraphLLM approach is shown in Figure 2. For more details on the scene graph representation, refer to Section 3.2. Our training process is two-stage. First, we pre-train the model on a dataset for various 3D scene understanding tasks using ground-truth instance segmentation. Next, we fine-tune 3DGraphLLM with predicted instance segmentation of scene point clouds, considering a scenario where ground-truth segmentation is unavailable (see Section 3.3).
Figure 2: The overall architecture of our approach. 3DGraphLLM leverages pre-trained encoders for 3D object point clouds and semantic relationships between objects. We introduce trainable layers to map the extracted graph node and edge features into the token embedding space of a pre-trained LLM. The scene graph is flattened for input into the LLM, with each object represented by a subgraph of its k nearest neighbors. To further adapt the LLM to 3D vision-language tasks, we add new object tokens to the LLM’s vocabulary and fine-tune it using LoRa.
@@ 3.1 Model Architecture
The model architecture includes pre-trained encoders for 3D point clouds and their semantic relationships, alongside a pre-trained LLM. We train projection layers to map the extracted object features and their relationships into the LLM’s token embedding space. Following the approach of Chat-Scene (Huang et al., 2024), we introduce additional object identifier tokens {<$OBJ+i>\}_{i=1}^{n}$ into the LLM’s vocabulary. Here and throughout, we use $n$ to denote the number of objects in the scene. These learned identifiers, along with the features from object subgraphs composed of nearest neighbors for each object, are used to create a flat representation of the scene graph, which is then fed into the LLM.
@@
Object Proposals. We use point clouds of objects in the scene as vertices in the scene graph $\mathit{G}$. In our experiments, we evaluate 3DGraphLLM in various modes, including ground-truth scene segmentation and instance segmentation using state-of-the-art neural network methods like Mask3D (Schult et al., 2023) and OneFormer3D (Kolodiazhnyi et al., 2024). Thus, the set $V$ of vertices of the graph consists of $n$ point clouds $\{P_{i}\}_{i=1}^{n}$, where $P_{i}\in\mathbb{R}^{m_{i}\times 6}$. Here, $m_{i}$ is the number of points in the $i$-th object proposal of instance segmentation of scene point cloud, and $6$ dimensions of each point correspond to its 3D coordinates and RGB color.
@@
Object Identifiers. Following the approach in Chat3D-Scene, we add a set of learnable identifier tokens {<$OBJ+i>\}_{i=1}^{n}$ to the LLM’s vocabulary for object identification. These tokens allow the model to identify objects in the scene by simply predicting the corresponding object identifier token. In our experiments, we assume a maximum of 200 objects per scene.
@@
2D Object Encoder. The results of Chat-Scene demonstrate that adding aggregated 2D DINOv2(Oquab et al., 2023) features increase the LLM performance on 3D vision-language tasks. Therefore, we add DINOv2 $Z_{i}^{2d}\in\mathbb{R}^{1\times 1024}$ features as an additional token describing the object subgraph. DINOv2 object features are obtained by aggregating features from the masked multi-view images where masks come from the projection of the object’s 3D point cloud.
@@
3D Object Encoder. We extract vertex features using a pre-trained Uni3D (Zhou et al., 2023b) encoder, which generates point cloud features aligned with their textual descriptions. Since this model is pre-trained on a large dataset, it enables us to produce high-quality graph vertex embeddings across various data domains. For each object point cloud $P_{i}$, we extract Uni3D feature $Z_{i}^{v_{p}}\in\mathbb{R}^{1\times 1024}$.
@@
Edge Feature Encoder. One challenge in generating features for semantic relationships between objects is that most methods for 3D semantic scene graph generation are trained on 3RScan scenes (Wald et al., 2019), while visual grounding tasks are typically tested on ScanNet scenes (Dai et al., 2017). Although both datasets belong to the indoor scene domain, existing methods struggle with performance in cross-domain testing, resulting in a drop in accuracy for the grounding task (Miyanishi et al., 2024).
@@
To extract semantic relationships between objects, we use VL-SAT (Wang et al., 2023a), a method for generating 3D semantic scene graphs from point clouds. One of its key advantages is that it only requires 3D point cloud coordinates as input during prediction, while leveraging knowledge transfer from the pre-trained CLIP model (Radford et al., 2021). This allows the method to perform well when applied to new scene domains (Wang et al., 2023a), as confirmed by our experiments (see Section 4.3 and Tables 3 and 4). For each pair of point clouds $P_{i}$ and $P_{j}$, we generate a latent feature representing their relationship $Z_{ij}^{e}\in\mathbb{R}^{1\times 512}$, which corresponds to VL-SAT graph neural network feature before the classification head assigning semantic categories to the graph edges. While VL-SAT predicts a fixed set of relationships between objects, these relationships are not mutually exclusive (e.g., "larger" and "close"). Therefore, we use latent features to capture possible combinations of these semantic relationships.
@@
2D/3D object, and semantic relation projection. To adapt the extracted features for the language model, we use three trainable projection modules: the 2D Object Projection $f_{2d}(\cdot)$, which maps the 2D image features of objects, the 3D Object Projection $f_{v}(\cdot)$, which maps the point cloud features of objects, and the Semantic Relation Projection $f_{e}(\cdot)$, which maps the features of semantic relationships between objects. Therefore, for the $i$-th object, the 2D and 3D object features are projected to token embeddings $F_{i}^{v}$ and $F_{i}^{2d}$ respectively. For the pair of $i$-th and $j$-th objects, the semantic relation feature is projected to token embedding $F_{ij}^{e}$:

### v3

@@ Abstract
A 3D scene graph represents a compact scene model by capturing both the objects present and the semantic relationships between them, making it a promising structure for robotic applications. To effectively interact with users, an embodied intelligent agent should be able to answer a wide range of natural language queries about the surrounding 3D environment. Large Language Models (LLMs) are beneficial solutions for user-robot interaction due to their natural language understanding and reasoning abilities. Recent methods for learning scene representations have shown that adapting these representations to the 3D world can significantly improve the quality of LLM responses. However, existing methods typically rely only on geometric information, such as object coordinates, and overlook the rich semantic relationships between objects. In this work, we propose 3DGraphLLM, a method for constructing a learnable representation of a 3D scene graph that explicitly incorporates semantic relationships. This representation is used as input to LLMs for performing 3D vision-language tasks. In our experiments on popular ScanRefer, Multi3DRefer, ScanQA, Sqa3D, and Scan2cap datasets, we demonstrate that our approach outperforms baselines that do not leverage semantic relationships between objects. The code is publicly available at https://github.com/CognitiveAISystems/3DGraphLLM.
@@ 1 Introduction
In this paper, we consider scene understanding in the context of 3D vision-language tasks: 3D referred object grounding task, 3D dense scene captioning and 3D visual question answering. The 3D referred object grounding task involves identifying a region within a 3D scene that corresponds to a natural language query. These queries often describe object properties (e.g., color, size) as well as spatial relationships (e.g., a mug on a table). A common setup of this problem assumes access to a 3D reconstruction of the scene, such as a point cloud, mesh, or NeRF. The objective is to predict the bounding box of the object or region referenced in the query. The goal of 3D dense scene captioning is to generate a textual description of a selected object in the 3D scene, including its attributes or relationships. Finally, the goal of the 3D visual question answering task is to generate text answers to various questions about the properties of the scene. It seems promising to explicitly use a three-dimensional scene graph to solve these tasks.
@@
Figure 1: The proposed 3DGraphLLM approach leverages 3D semantic scene graph learnable representation supplied as input to an LLM to perform various 3D vision-language tasks.
A 3D scene graph provides a unified representation of a scene by storing multimodal information about individual objects, along with their semantic relationships [52, 32] and hierarchical organization [53, 20]. It also supports real-time updates in dynamic environments, making it suitable for interactive scenes [45, 38]. Furthermore, representing the scene as a graph enables the use of graph algorithms for tasks such as navigation [64, 19, 20] and object search based on textual queries [14, 4, 53, 17].
Solving 3D vision-language tasks is essential for embodied intelligent agents [5, 9, 3]. To interact effectively with users, such agents must be able to describe their environment and answer questions about its properties using natural language. Large Language Models (LLMs) are particularly well-suited for this, thanks to their strong capabilities in language understanding and commonsense reasoning. They can interpret user queries and match them to objects in a scene, even when the queries are vague or indirect [22, 51, 17]. By leveraging LLMs, it becomes easier to adapt the method to new object categories and relationships mentioned in referring expressions. LLMs can also handle complex queries that describe an object by its function rather than its name (e.g., ”somewhere to sit”).
A 3D scene can be represented for input to an LLM either as text [17, 34, 53, 20, 55, 58] or as a implicit learnable representation [22, 7, 24, 8, 10]. Learnable representations encode objects and their relationships into embeddings, using significantly fewer tokens than textual descriptions. This compact form not only increases the speed of LLM inference but also enhances response quality by enabling better adaptation to 3D scenes. However, existing methods [22, 7, 24, 8] that use learnable 3D scene representations for vision-language tasks typically rely only on spatial coordinates and fail to incorporate semantic relationships between objects - limiting the expressiveness and reasoning capabilities of the model.
@@
In this paper, we introduce 3DGraphLLM, a novel learnable representation of a 3D scene graph designed for use as input to an LLM (see  Fig. 1). The representation consists of a list of learnable embeddings for scene objects, where each object is modeled as a local subgraph that includes the object itself and its nearest neighbors. These subgraphs are provided to the LLM as a sequence of triplets (object1, relation, object2). Semantic relations are encoded using features derived from the semantic edges of the scene graph, generated by state-of-the-art methods such as VL-SAT [52]. Our experiments show that incorporating semantic relationships between objects significantly improves the accuracy of LLM responses in 3D vision-language tasks, outperforming baseline methods that use learnable scene representations without semantic context.
@@
  • We introduce 3DGraphLLM, the first method for creating a learnable 3D scene graph representation specifically designed for LLMs. It enables semantic relationships between objects in a scene to be mapped directly into the LLM’s token embedding space.
  • We propose an algorithm that generates a flat sequence of graph embedding tokens by selecting object subgraphs using k-nearest neighbors with Non-Maximum Supression (NMS) and a minimum distance filters between objects. This approach reduces the number of tokens needed to describe the scene, thereby improving inference speed.
  • 3DGraphLLM outperforms the baseline method which does not use semantic relationships on the 3D referred object grounding task, achieving improvements of +7.5% F1@0.5 on the Multi3DRefer[60] and +6.4% Acc@0.5 on ScanRefer [5] benchmarks. It also improves performance on 3D scene captioning, with a +3.9% CIDEr@0.5 score on the Scan2Cap [9] dataset. 3DGraphLLM achieves state-of-the-art results in 3D referred object grounding while requiring up to five times less inference time compared to LVLM-based methods.
@@ 2 Related works
3D Language Scene Understanding. 3D scene understanding is a complex computer vision task that involves identifying the semantic, physical, and functional properties of objects, as well as their mutual relations. One of the goals of 3D scene understanding is to develop methods capable of responding to natural language queries about the scene. The queries may correspond to different visual-language tasks such as 3D referred object grounding [5, 60, 36], question answering [3], and dense scene captioning [9]. Recent approaches address these queries by reconstructing the scene as a 3D mesh [41] or point cloud [61, 6, 65], often enhanced with instance segmentation [65].
@@
The emergence of transformer models [48] has enabled the development of neural network models that create a learnable representation of a scene for answering various language queries. MultiCLIP [12] proposes to align 3D scene representation with text queries and multi-view 2D CLIP [44] embeddings to improve the quality of question answering. 3DVG-Transformer [61] and Vil3DRef [6] methods introduce modules for modeling spatial relationships between objects to improve the quality of object grounding. 3D-VisTA [65] presents a transformer model for aligning 3D object and text representations, coupled with an unsupervised pre-training scheme to solve various 3D vision-text problems using specialized task-specific heads. However, these approaches face challenges in generalizing to new tasks and domains. In contrast, leveraging large language models (LLMs) for scene understanding enhances generalization capabilities and taps into the extensive knowledge LLMs contain about the physical world [22].
Scene Graphs. The concept of a scene graph was initially developed for 2D images, providing a structured representation of a scene’s semantics by incorporating relationships between the semantic elements [29]. In the context of images, scene graphs have proven effective for tasks such as content-based image retrieval [29, 40], 2D referring expression comprehension [56, 47, 18], image caption [57, 42], image generation [30, 13].
In 3D scenes, a scene graph is commonly used to address robotics challenges such as planning [53, 20], object grounding for navigation [53, 17, 34, 20] and manipulation [20], as well as scene generation [59, 16]. Our approach is part of a class of methods that utilize an implicit representation of the scene graph, such as OVSG [4], which frames the problem of 3D object grounding as subgraph retrieval. 3DGraphQA [54] proposes to use the bilinear graph neural network for feature fusion between scene and question graphs for question answering task. FFL-3DOG [14] builds a graph based on a text query, which is used to refine the visual graph to select from its vertices the one that best fits the description. However, the application scope of this method is limited to specific tasks such as 3D referred object grounding or question answering. In contrast, we propose a more versatile method capable of solving various 3D vision-language tasks.
@@
Large Language Models for Scene Understanding. Large language models (LLMs) offer several advantages for scene understanding, notably enhancing the ability to address complex queries that require common knowledge. LLMs can serve as agents that decompose user queries into elementary tasks, which can then be addressed by other methods [55, 58]. Additionally, LLMs can act as an interface for reasoning by processing textual descriptions of the scene as input [34, 17]. BBQ [34] and ConceptGraphs [17] demonstrate that using a text-based graph representation with an LLM interface significantly improves the quality of object retrieval compared to using CLIP features of objects. HOV-SG [53] constructs a hierarchical graph consisting of objects, rooms, and floors, and demonstrates the effectiveness of such a representation for the task of object grounding given a query containing object location hints. The authors of the MOMA [20] method propose using a hierarchical scene graph together with a navigational Voronoi graph as input to LLM to predict a high-level policy for object search for navigation and manipulation. However, using text to describe an object in a scene graph inevitably leads to the loss of some of the information contained in its RGB point cloud. Additionally, in the case of using a text graph, several hundred tokens may be required to describe one object (its semantic class, pose), which will significantly slow down LLM inference in the case of a large number of objects in the scene.
@@
Recent advancements have successfully integrated point cloud data into LLMs by employing pre-trained point cloud encoders and training adapters to align the resulting representations with the LLM embedding space. 3D-LLM [21] aggregates 3D point cloud features from a sequence of 2D images and then solves the grounding problem as a prediction of a sequence of location tokens added to the LLM dictionary. Chat-Scene [25] generates 2D and 3D features for each object in the scene and introduces learnable object identifier tokens to solve object grounding, dense scene captioning, and question answering problems. LLA3D [7] proposes to use a set of trainable fixed-length query tokens obtained by interacting potential visual cues, text cues, and object point cloud features in a transformer model. Grounded 3D-LLM [8] uses referent tokens to decode object masks in point clouds. Additionally, research has demonstrated that incorporating spatial information, such as object coordinates [24] or depth maps [10], enhances the accuracy of responses to user queries.
@@ 3 Method
Our approach uses a set of point clouds of scene objects as input. The objects’ point clouds can be obtained either from ground-truth annotations or through state-of-the-art point cloud instance segmentation methods. These point clouds are used to extract scene graph features (see  Sec. 3.1). A scene graph consists of nodes representing the objects and edges corresponding to semantic relationships between them. To convert the scene graph into a token sequence, we represent each object by an identifier, its 2D object feature, and a subgraph comprising the object’s k nearest neighbors. The relationships between an object and its neighbors are encoded as triplets $(object_{i},relation_{ij},object_{j})$. The scheme of the 3DGraphLLM approach is shown in  Fig. 2. For more details on the scene graph representation, refer to  Sec. 3.2. Our training process is two-stage. First, we pre-train the model on a dataset for various 3D scene understanding tasks using ground-truth instance segmentation. Next, we fine-tune 3DGraphLLM with predicted instance segmentation of scene point clouds, considering a scenario where ground-truth segmentation is unavailable (see  Sec. 3.3).
Figure 2: The overall architecture of our approach. We introduce trainable layers to map the extracted graph node and edge features into the token embedding space of a pre-trained LLM. The scene graph is flattened for input into the LLM, with each object represented by a subgraph of its k nearest neighbors. To further adapt the LLM to 3D vision-language tasks, we add new object tokens to the LLM’s vocabulary alongside with objects’ 2D features and fine-tune the LLM using LoRa.
@@ 3.1 Model Architecture
The model architecture includes pre-trained encoders for 2D images, 3D point clouds, and point clouds semantic relationships, alongside a pre-trained LLM. We train projection layers to map the extracted object features and their relationships into the LLM’s token embedding space. Following the approach of Chat-Scene [25], we introduce additional object identifier tokens $\{<\verb+OBJ+i>\}_{i=1}^{n}$ into the LLM’s vocabulary. Here and throughout, we use $n$ to denote the number of objects in the scene. These learned identifiers, with the features from object subgraphs composed of nearest neighbors for each object, are used to create a flat representation of the scene graph, which is then fed into the LLM.
@@
Object Proposals. We use point clouds of objects in the scene as vertices in the scene graph $\mathit{G}$. In our experiments, we evaluate 3DGraphLLM in various modes, including ground-truth scene segmentation and instance segmentation using state-of-the-art neural network methods like Mask3D [46] and OneFormer3D [33]. Thus, the set $V$ of vertices of the graph consists of $n$ point clouds $\{P_{i}\}_{i=1}^{n}$, where $P_{i}\in\mathbb{R}^{m_{i}\times 6}$. Here, $m_{i}$ is the number of points in the $i$-th object proposal of instance segmentation of scene point cloud, and $6$ dimensions of each point correspond to its 3D coordinates and RGB color.
@@
Object Identifiers. Following the approach in Chat-Scene, we add a set of learnable identifier tokens $\{<\verb+OBJ+i>\}_{i=1}^{n}$ to the LLM’s vocabulary for object identification. These tokens allow the model to identify objects in the scene by simply predicting the corresponding object identifier token. In our experiments, we assume a maximum of 200 objects per scene.
@@
2D Object Encoder. The results of Chat-Scene demonstrate that adding aggregated 2D DINOv2[37] features increases the LLM performance on 3D vision-language tasks. Therefore, we add DINOv2 $Z_{i}^{2d}\in\mathbb{R}^{1\times 1024}$ features as an additional token describing the object subgraph. DINOv2 object features are obtained by aggregating features from the masked multi-view images where masks come from the projection of the object’s 3D point cloud.
@@
3D Object Encoder. We extract vertex features using a pre-trained Uni3D [63] encoder, which generates point cloud features aligned with their textual descriptions. Since this model is pre-trained on a large dataset, it enables us to produce high-quality graph vertex embeddings across various data domains. For each object point cloud $P_{i}$, we extract Uni3D feature $Z_{i}^{v_{p}}\in\mathbb{R}^{1\times 1024}$.
@@
Edge Feature Encoder. One challenge in generating features for semantic relationships between objects is that most methods for 3D semantic scene graph generation are trained on 3RScan scenes [50], while visual grounding tasks are typically tested on ScanNet scenes [11]. Although both datasets belong to the indoor scene domain, existing methods struggle with performance in cross-domain testing, resulting in a drop in accuracy for the grounding task [36].
@@
To extract semantic relationships between objects, we use VL-SAT [52], a method for generating 3D semantic scene graphs from point clouds. One of its key advantages is that it only requires 3D point cloud coordinates as input during prediction while leveraging knowledge transfer from the pre-trained CLIP model [44]. This allows the method to perform well when applied to new scene domains [52], as confirmed by our experiments (see  Sec. 4.2). For each pair of point clouds $P_{i}$ and $P_{j}$, we generate a latent feature representing their relationship $Z_{ij}^{e}\in\mathbb{R}^{1\times 512}$, which corresponds to VL-SAT graph neural network feature before the classification head assigning semantic categories to the graph edges. While VL-SAT predicts a fixed set of relationships between objects, these relationships are not mutually exclusive (e.g., ”larger” and ”close”). Therefore, we use latent features to capture possible combinations of these semantic relationships.
@@
2D/3D object, and semantic relation projection. To adapt the extracted features for the language model, we use three trainable projection modules: the 2D Object Projection $f_{2d}(\cdot)$, which maps the 2D image features of objects, the 3D Object Projection $f_{v}(\cdot)$, which maps the point cloud features of objects, and the Semantic Relation Projection $f_{e}(\cdot)$, which maps the features of semantic relationships between objects. Therefore, for the $i$-th object, the 2D and 3D object features are projected to token embeddings $F_{i}^{v}$ and $F_{i}^{2d}$, respectively. For the pair of $i$-th and $j$-th objects, the semantic relation feature is projected to token embedding $F_{ij}^{e}$: